OpenStudy (anonymous):

The weight of water 62 1/2 lb per cubic foot. What is the weight of 29 1/8 cubic feet of water? Simplify the answer using whole number, proper fraction, or mixed number.

OpenStudy (ash2326):

1 cubic foot of water weighs 62 1/2 lb or 125/2 pounds So the weight of 29 1/8 or 233/8 cubic foot water is = 233/8* 125/2= 1820.3125 lbs or 1820 5/16 lbs

OpenStudy (anonymous):

Thank you.